# AnomalyDINO: Boosting Patch-based Few-shot Anomaly Detection with DINOv2

### v1

@@ C.3 Architecture Size
DINOv2 is available in different distillation sizes (S, B, L, and G). This section analyzes the implications of choosing different backbone sizes for AnomalyDINO. The comparison including the best competing methods is depicted in Figure 12.
We see that different architecture sizes have some effect on the image-level AUROC. AnomalyDINO-L outperforms smaller architectures, S (our default) and B, on VisA in the few-shot setting with $k\in\{1,2,4\}$. For $k\in{8,16}$ AnomalyDINO-L is on par with AnomalyDINO-S (and APRIL-GAN). For $k=16$ all architecture sizes outperform the best competitor (APRIL-GAN).
On MVTec-AD, we see that AnomalyDINO-S performs slightly better, but the gap between AnomalyDINO-S and AnomalyDINO-B reduces with increasing shot size. In particular, all architecture sizes outperform the closest competitor (ADP).
Refer to captionRefer to caption
Figure 12: Effect of model size on detection AUROC. We report mean and standard deviation for the 1-shot setting on MVTec-AD (left) and VisA (right) over three seeds. Note, that the results for WinCLIP+ for 8 and 16 shots are those of the WinCLIP re-implementation by Kwak et al. (2024).
@@ Appendix D Extending AnomalyDINO to the Batched Zero-Shot Setting
We extend the proposed method to the batched zero-shot setting. Recall, that in this setting all test samples $X_{\mathrm{test}}$ are provided (or at least a sufficiently large batch), but no (labeled) training or reference samples. The underlying (and necessary) assumption to meaningfully predict anomalies solely based on test samples, is that the majority of samples (or in our case, patches) at test time are from the nominal data distribution (see e.g., Li et al., 2024a).
We need to alter our method, outlined in Section 3, only slightly to adapt it to the batched setting. We score a test sample $\mathbf{x}^{(j)}\in X_{\mathrm{test}}$ in comparison to all remaining test sample $X_{\operatorname{test}}\setminus\{\mathbf{x}^{(j)}\}$, following the idea of mutual scoring (Li et al., 2024b). For each test sample $\mathbf{x}^{(j)}\in X_{\mathrm{test}}$ we therefore collect all patch representations not belonging to $\mathbf{x}^{(j)}$ in a memory bank, again utilizing DINOv2 as patch-level feature extractor $f$,
$\mathcal{M}_{j}:=\bigcup_{\mathbf{x}^{(i)}\in X_{\operatorname{test}}\setminus\{\mathbf{x}^{(j)}\}}\bigl{\{}\mathbf{p}_{m}\;|\;f(\mathbf{x}^{(i)})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n}),\,m\in\{1,\dots,n\}\bigr{\}}\enspace.$ (5)
@@
where $L_{\alpha}(\mathcal{D})$ contains the values below the $\alpha$ quantile of $\mathcal{D}$. We set $\alpha=0.1\%$ as anomalous patches are rare by assumption, and $\mathcal{D}(\mathbf{p}_{\mathrm{test}},\mathcal{M}_{j})$ large enough to accurately estimate the tail of the distribution. [^6] For MVTec-AD the total number of test patches extracted by DINOv2 (ViT-S) range between 43.008 and 171.008 per category, and for VisA between 163.200 and 334.464. The image-level score $s(\mathbf{x}_{\mathrm{test}})$ is again given by aggregating the patch-level anomaly score following Equation 4. Computation of the cosine distances and the proposed aggregation statistics can be effectively implemented as matrix operations on GPU such that the batched zero-shot inference time for AnomalyDINO amounts to roughly 0.1 seconds/sample for MVTec-AD (again measured on an NVIDIA A40; feature extraction takes approximately the same amount of time).
Some resulting anomaly maps in the batched zero-shot setting are depicted in Figure 13.
Figure 13: Anomaly maps for the batched zero-shot setting on MVTec-AD. The left-most sample in each category is a ‘good’ test sample for reference, followed by four samples with anomalies.

### v2

@@ C.3 Architecture Size
DINOv2 is available in different distillation sizes (S, B, L, and G). This section analyzes the implications of choosing different backbone sizes for AnomalyDINO. The comparison including the best competing methods is depicted in Figure 14. We see that different architecture sizes have indeed an effect on the image-level AUROC.
On MVTec-AD, we see that AnomalyDINO-S performs best, followed by the next larger model AnomalyDINO-B (which might contrast the common belief that larger models always perform better). In particular, all architecture sizes outperform the closest competitor (ADP).
Regarding VisA, larger architectures slightly outperform our default setting, AnomalyDINO-S. All architecture sizes are on par with APRIL-GAN at $k=4$, but outperform all competitors for $k>4$. Figure 14 also showcases that the performance of the proposed method scales preferably with the number of reference samples.
(a) Detection ARUOC for MVTec-AD.
(b) Detection ARUOC for VisA.
Figure 14: Effect of model size on detection AUROC for MVTec-AD and VisA. We report mean and standard deviation for the 1-shot setting over three seeds. The image resolution of AnomalyDINO is set to 672. Note, that the results for WinCLIP+ for 8 and 16 shots are those of the WinCLIP re-implementation by [20].
@@ Appendix D Extending AnomalyDINO to the Batched Zero-Shot Setting
We extend the proposed method to the batched zero-shot setting. Recall, that in this setting all test samples $X_{\mathrm{test}}$ are provided (or at least a sufficiently large batch), but no (labeled) training or reference samples. The underlying (and necessary) assumption to meaningfully predict anomalies solely based on test samples, is that the majority of samples (or in our case, patches) at test time are from the nominal data distribution (see e.g., [21]).
We need to alter our method, outlined in Section 3, only slightly to adapt it to the batched setting. We score a test sample $\mathbf{x}^{(j)}\in X_{\mathrm{test}}$ in comparison to all remaining test sample $X_{\operatorname{test}}\setminus\{\mathbf{x}^{(j)}\}$, following the idea of mutual scoring [23]. For each test sample $\mathbf{x}^{(j)}\in X_{\mathrm{test}}$ we therefore collect all patch representations not belonging to $\mathbf{x}^{(j)}$ in a memory bank, again utilizing DINOv2 as patch-level feature extractor $f$,
$\mathcal{M}_{j}:=\bigcup_{\mathbf{x}^{(i)}\in X_{\operatorname{test}}\setminus\{\mathbf{x}^{(j)}\}}\bigl{\{}\mathbf{p}_{m}\mid f(\mathbf{x}^{(i)})=(\mathbf{p}_{1},\dots,\mathbf{p}_{n}),\,m\in[n]\bigr{\}}\enspace.$ (5)
@@
where $L_{\alpha}(\mathcal{D})$ contains the values below the $\alpha$ quantile of $\mathcal{D}$. We set $\alpha=0.1\%$ as anomalous patches are rare by assumption, and $\mathcal{D}(\mathbf{p}_{\mathrm{test}},\mathcal{M}_{j})$ large enough to accurately estimate the tail of the distribution. [^6] For MVTec-AD the total number of test patches extracted by DINOv2 (ViT-S) at a resolution of 448 range between 43.008 and 171.008 per category, and for VisA between 163.200 and 334.464. The image-level score $s(\mathbf{x}_{\mathrm{test}})$ is again given by aggregating the patch-level anomaly score following Equation 4. Computation of the cosine distances and the proposed aggregation statistics can be effectively implemented as matrix operations on GPU such that the batched zero-shot inference time for AnomalyDINO amounts to roughly 60 ms/sample for MVTec-AD at a resolution of 448 (again measured on an NVIDIA A40 GPU). Some resulting anomaly maps in the batched zero-shot setting are depicted in Figure 15.
Figure 15: Anomaly maps for the batched zero-shot setting on MVTec-AD. The left-most sample in each category is a ‘good’ test sample for reference, followed by four randomly picked samples with anomalies.